OpenStudy (anonymous):

MEdal x+7=6x-3 a. –5 b. –15 c. 2 d. –10 I thought it was -15 idk if im right tho

OpenStudy (anonymous):

x+7=6x-3 -x -x 7=5x-3 +3 +3 10=5x /5 /5 x=2 final answer: x=2